THE CHLORINE NUCLEAR QUADRUPOLE COUPLING CONSTANT IN HSiCl

\underline{WEI LIN} and STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459; MASARU FUKUSHIMA^{a}^{b} and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB, Canada, T6G 2G2.

HSiCl was produced by a pulsed discharge of free expansions of (CH_{3})_{3}SiCl or HSiCl_{3} diluted in a rare gas carrier. The 1_{01} - 0_{00} microwave transitions in H^{28}Si^{35}Cl, H^{28}Si^{37}Cl, H^{29}Si^{35}Cl, H^{30}Si^{35}Cl were measured at 14 - 15 GHz by FTMW spectroscopy. The effective rotational constants were obtained and the chlorine nuclear quadrupole coupling constants of the four isotopomers were determined. These hyperfine constants will be discussed and compared with those of related chlorosilylenes and chlorosilanes.

aPermanent address: Faculty of Information Sciences, Hiroshima City University, Asa-Minami, Hiroshima 731-3194,Japan

bSupport by Hiroshima City University President-Designated Grant for Long-Term Research Abroad for Young Researchers is gratefully acknowledged.